Could you please help me solve this problem?
Number problems. If the square of 2 less than an integer is 16, find the number(s).
Thank you,

Answer by checkley71(8403)

N-2^2=16
N+4=16
N=16-4
N=12 ANSWER.
PROOF
12+4=16
16=16